# Finite and full scale localization for the multi-frequency quasi-periodic CMV matrices

### v2

@@ Abstract
This paper formulates the finite and full-scale localization for multi-frequency quasi-periodic CMV matrices. This can be viewed as the CMV counterpart to the localization results by Goldstein, Schlag, and Voda [15, 16] on multi-frequency quasi-periodic Schrödinger operators.
@@
CMV matrix; Multi-frequency; Anderson localization; No double resonances condition; Factorization of characteristic determinant.
MSC:
[2010] 37A30 , 42C05 , 70G60
journal: aXiv
@@ 1 Introduction
Anderson localization (AL) describes the transformation of an impurity doped medium from a conductor to an insulator and this interesting phenomenon was discovered in physical experiments; see the Anderson’s seminal paper [1]. However, AL also holds significant mathematical importance. Mathematically, it implies that the corresponding operator has pure point spectrum with exponentially decaying eigenfunctions. We refer the reader to [11, 22] for the history of AL. Here we would like to mention that AL problems for Schrödinger operators are widely studied and fruitful results have been obtained, especially for quasi-periodic potentials with single frequency; see for example, [2, 4, 5, 8, 11, 12, 23] and references therein.
@@
Given the development of AL theory for Schrödinger operators, it is natural to investigate the Schrödinger operators with multi-frequency quasi-periodic (MF-QP) potentials. But it is a consensus that shifts on a multidimensional torus $\mathbb{T}^{d}$ turned out to be harder to analyze and the theory was less developed. Chulaevsky and Sinai [10] originally established AL for two-frequency Schrödinger operator almost everywhere in phase space. For the frequency of more than two dimensional case, the same result was obtained in [4]. However, these results were established directly in infinite volume, whereas localization in finite volume was not studied. Recently, Goldstein, Schlag and Voda [15, 16] initially developed some analytical techniques and established both the finite and full scale localization of the MF-QP Schrödinger operators. Based on the tools obtained in [15], Goldstein, Schlag and Voda [16] were committed to a class of multi-frequency quasi-periodic Schrödinger operators and they derived a surprising result that the spectrum of this kind of operators is actually a single interval. This proves a conjecture proposed by Chulaevsky and Sinai in [10]. Here we would also like to mention the recent papers [28] and [29] on MF-QP operators.
As for CMV matrices, substantial progress has been made in developing the mathematical theory of localization. In [9], Cedzich and Werner studied the one-dimensional electric quantum walks, they found out that the absolutely continuous spectrum of the quantum walk with a kind of coins is empty. In addition, as a result of the connection between one-dimensional quantum walks and CMV matrices, the AL result for the walk implies AL for the CMV matrices with skew-shift Verblunsky coefficients. In [20], Lin, Guo and Piao extended the results of Bourgain, Goldstein, Schlag [6] and derived the positivity of the Lyapunov exponents and AL for a class of CMV matrices whose Verblunsky coefficients are defined by skew-shift. Wang and Damanik [27] extended the results by Bourgain and Schlag [5] and obtained AL for the quasi-periodic CMV matrices with single frequency. Zhu [30] established both AL and strong dynamical localization for random CMV matrices. Lin, Guo and Piao [21] proved AL for CMV matrices with Verblunsky coefficients generated by hyperbolic toral automorphism. Recently, for CMV matrices with small and quasi-periodic Verblunsky coefficients, Li, Damanik and Zhou [18] proved that the associated spectral measure is purely absolutely continuous on the essential spectrum. In another paper [19], they established Cantor spectrum for some kind of MF-QP CMV matries with Verblunsky coefficients distributed on a circle inside $\partial\mathbb{D}$ by the Gap Labeling Theorem. However, we still do not really understand much about the MF-QP CMV matrices.
@@
To our knowledge, no research paper has addressed the finite and full-scale localization for MF-QP CMV matrices. This paper aims to fill this gap in the literature. To do that, we need large deviation theorem (LDT) for the entries of the transfer matrix. Comparing with the results in [13], there are some difficulties we have to overcome. This is primarily due to the more complex structure of CMV matrices. It is known to all, Schrödinger operator is a real tri-diagonal matrix, while the CMV matrix is a complex five-diagonal matrix, then the analysis about the corresponding characteristic determinant (entries of the transfer matrix) is more complicated. To treat this issue, we decompose the matrix into two parts and consider two cases; see the proof of Lemma 3.15 in detail. Consequently, the large deviation theorem for the entries of the transfer matrix follows.
@@
In addition, considering the restriction of the extended CMV matrix to a finite interval, from [17] and [30], the Poisson formula of CMV matrices depends on the parity of the endpoints of this interval. This also makes the analysis of the covering form of LDT more difficult to some extent. Thus we consider a subinterval of the above finite interval instead, then we get the covering lemma. Since the subinterval is a closed interval, the covering lemma can be applied to the original interval. And the covering form of LDT is obtained.
@@
Furthermore, the restriction of the Schrödinger operator to a finite interval results in a Hermitian matrix with many favorable properties. In contrast, the restriction of the CMV matrix results in a unitary matrix after modifying the boundary conditions, and its properties are less favorable than those of a Hermitian matrix. For example, there is a famous Cauchy Interlacing Theorem for the eigenvalues of Hermitian matrices. But we can not expect such a result for unitary matrices. This directly results in a challenge in estimating the number of zeros for the characteristic determinant within a disk. Thus, the factorization of the characteristic determinant becomes more complex.
In contrast to the results on localization for quasi-periodic CMV matrices with single frequency [27], we provide a more detailed description of the elimination of double resonances. In [27], one just needs to eliminate a set of frequency satisfying double resonances condition by fixing some phase. While in our paper, we find the sets of frequency and phase respectively, then we prove the full scale localization for all frequencies and phases outside these two sets. In contrast to the results on localization for quasi-periodic CMV matrices with single frequency[27], we provide a more detailed description of the elimination of double resonances. Naturally, the process becomes more complex.
@@
The rest of this paper is organized as follows. We describe the settings in Section 2. In Section 3, we give the estimates of the Lyapunov exponent and prove LDT for the transfer matrix and the entries of the transfer matrix respectively. Furthermore, with the aid of Cartan’s estimate and Poisson formula, we get the spectral and covering form of LDT. Based on the results derived in Section 3, we prove the finite scale localization of the MF-QP CMV matrices in Section 4, see Theorem 3.27. We introduce the no double resonances (NDR) condition in Section 5. We study the factorization of the characteristic determinant under NDR condition in Section 6. In Section 7, we derive a result on elimination of double resonances by applying the theory of semialgebraic sets. In Section 8, we study the elimination of double resonances between NDR intervals. Based on the analysis in previous sections, we obtain the full scale localization of the MF-QP CMV matrices in Section 9, see Theorem 6.7. We put some useful definitions and lemmas in the appendix section, that is, Section 10.
@@ 2 Preliminaries
Let $\mathbb{D}=\{z\mid|z|<1\}$ denote the open unit disk in $\mathbb{C}$, and let $\mu$ be a nontrivial probability measure on $\partial\mathbb{D}=\{z\mid|z|=1\}$. Consequently, the functions $1,z,z^{2},\ldots$ are linearly independent in the Hilbert space $L^{2}(\partial\mathbb{D},d\mu)$. Let $\Phi_{n}(z)$ be the monic orthogonal polynomials, that is, $\Phi_{n}(z)=P_{n}[z^{n}]$, $P_{n}\equiv$ projection onto $\{1,z,z^{2},\ldots,z^{n-1}\}^{\bot}$ in $L^{2}(\partial\mathbb{D},d\mu)$. The orthonormal polynomials are then given by
@@
Then the Verblunsky coefficients $\{\alpha_{n}\}_{n=0}^{\infty}$ ($\alpha_{n}\in\mathbb{D}$) obey the equation
@@
Applying the Gram-Schmidt procedure to $1,z,z^{-1},z^{2},z^{-2},\cdots$, we obtain a CMV basis $\{\chi_{j}\}_{j=0}^{\infty}$ of $L^{2}(\partial\mathbb{D},d\mu)$, and the matrix representation of multiplication by $z$ relative to the CMV basis gives rise to the CMV matrix $\mathcal{C}$,
@@
$\mathcal{C}=\left[\begin{array}[]{ccccccc}\overline{\alpha}_{0}&\overline{\alpha}_{1}\rho_{0}&\rho_{1}\rho_{0}&&&&\\ \rho_{0}&-\overline{\alpha}_{1}\alpha_{0}&-\rho_{1}\alpha_{0}&&&&\\ &\overline{\alpha}_{2}\rho_{1}&-\overline{\alpha}_{2}\alpha_{1}&\overline{\alpha}_{3}\rho_{2}&\rho_{3}\rho_{2}&&\\ &\rho_{2}\rho_{1}&-\rho_{2}\alpha_{1}&-\overline{\alpha}_{3}\alpha_{2}&-\rho_{3}\alpha_{2}&&\\ &&&\overline{\alpha}_{4}\rho_{3}&-\overline{\alpha}_{4}\alpha_{3}&\overline{\alpha}_{5}\rho_{4}&\\ &&&\rho_{4}\rho_{3}&-\rho_{4}\alpha_{3}&-\overline{\alpha}_{5}\alpha_{4}&\\ &&&&\ddots&\ddots&\ddots\end{array}\right],$
@@
$\mathcal{E}=\left[\begin{array}[]{cccccccc}\ddots&\ddots&\ddots&&&&&\\ &-\overline{\alpha}_{0}\alpha_{-1}&\overline{\alpha}_{1}\rho_{0}&\rho_{1}\rho_{0}&&&&\\ &-\rho_{0}\alpha_{-1}&-\overline{\alpha}_{1}\alpha_{0}&-\rho_{1}\alpha_{0}&&&&\\ &&\overline{\alpha}_{2}\rho_{1}&-\overline{\alpha}_{2}\alpha_{1}&\overline{\alpha}_{3}\rho_{2}&\rho_{3}\rho_{2}&&\\ &&\rho_{2}\rho_{1}&-\rho_{2}\alpha_{1}&-\overline{\alpha}_{3}\alpha_{2}&-\rho_{3}\alpha_{2}&&\\ &&&&\overline{\alpha}_{4}\rho_{3}&-\overline{\alpha}_{4}\alpha_{3}&\overline{\alpha}_{5}\rho_{4}&\\ &&&&\rho_{4}\rho_{3}&-\rho_{4}\alpha_{3}&-\overline{\alpha}_{5}\alpha_{4}&\\ &&&&&\ddots&\ddots&\ddots\end{array}\right].$
In this paper, we consider a sequence of Verblunsky coefficients generated by an analytic function $\alpha(\cdot):\mathbb{T}^{d}\rightarrow\mathbb{D}$, $\alpha_{n}(x)=\alpha(T^{n}x)=\alpha(x+n\omega)$, $d>1$. Here, $T$ is the shift map defined by $Tx=x+\omega$, $\mathbb{T}:=\mathbb{R}/\mathbb{Z}$, and $x,\omega\in\mathbb{T}^{d}$ are referred to as phase and frequency respectively. Moreover, the frequency vector $\omega\in\mathbb{T}^{d}$ satisfies the standard Diophantine condition
@@
[…]
@@
Applying the Szegő dual to both sides of (2.5), we obtain
@@
$\left[\begin{array}[]{c}\varphi_{n+1}\\ \varphi_{n+1}^{*}\end{array}\right]=S(\omega,z;x+n\omega)\left[\begin{array}[]{c}\varphi_{n}\\ \varphi_{n}^{*}\end{array}\right],$
@@
$S(\omega,z;x)=\frac{1}{\rho(x)}\left[\begin{array}[]{cc}z&-\overline{\alpha}(x)\\ -\alpha(x)z&1\end{array}\right].$
@@
$M(\omega,z;x)=\frac{1}{\rho(x)}\left[\begin{array}[]{cc}\sqrt{z}&-\frac{\overline{\alpha}(x)}{\sqrt{z}}\\ -\alpha(x)\sqrt{z}&\frac{1}{\sqrt{z}}\end{array}\right]\in\mathbb{S}\mathbb{U}(1,1),$
@@
$\lim_{n\rightarrow\infty}u_{n}(\omega,z;x)=\lim_{n\rightarrow\infty}L_{n}(\omega,z)=L(\omega,z).$ (2.12)
exists for $a.e.\;x$. Note that
@@ 3.1 LDT Estimate for the Monodromy
[14, Proposition 9.1] Let $d$ be a positive integer. Suppose $u:D(0,2)^{d}\rightarrow[-1,1]$ is subharmonic in each variable; i.e., $z\rightarrow u(z_{1},u_{2},\ldots,z_{d})$ is subharmonic for any choice of $(z_{2},\ldots,z_{d})\in D(0,2)^{d-1}$ and similarly for each of the other variables. Assume furthermore that for some $n\geq 1$
@@
[…]
Based on the above lemma, obtain the following LDT, a fundamental tool in localization theory.
@@
Then $u_{n}$ is a continuous subharmonic function bounded by $1$ in $D(0,1)^{d}$. Moreover, $u_{n}$ satisfies the conditions in the above lemma, and (3.3) follows directly from (3.2). ∎

### v3

@@ Abstract
This paper establishes finite and full scale localization for multi-frequency quasi-periodic CMV matrices, filling a key gap as the CMV counterpart to Goldstein-Schlag-Voda’s multi-frequency quasi-periodic Schrödinger operator results [15, 16]. The multi-frequency case presents significant additional challenges due to the complex five-diagonal structure of CMV matrices, the unitary (rather than Hermitian) nature of their finite-volume restrictions. We overcome these obstacles by proving a large deviation theorem for the entries of the transfer matrix, deriving a covering form of the large deviation theorem applicable to CMV matrices. Furthermore, we provide a detailed treatment of double resonance elimination via the theory of semialgebraic sets and no double resonances condition, which is more intricate than in the single-frequency case.
@@
CMV matrix; Anderson localization; Multi-frequency quasi-periodic potentials; Lyapunov exponents; Spectral theory.
2020 MSC:
37A05 , 42C05 , 70G60
journal: arXiv
@@ 1 Introduction
Anderson localization (AL), originating from Anderson’s seminal work on disordered systems [1], describes the transition of a medium from a conductor to an insulator. Mathematically, it manifests as the operator possessing pure point spectrum with exponentially decaying eigenfunctions, a property of profound mathematical significance [11, 22]. Here we would like to mention that AL problems for Schrödinger operators are widely studied and fruitful results have been obtained, especially for quasi-periodic potentials with single-frequency; see for example, [2, 4, 5, 8, 11, 12, 23] and references therein.
@@
Given the well-developed AL theory for single-frequency quasi-periodic Schrödinger operators, extending it to the multi-frequency case is a natural yet challenging direction. The analysis of shifts on a higher-dimensional torus $\mathbb{T}^{d}$ introduces significant complexities, resulting in a less developed theory for multi-frequency quasi-periodic (MF-QP) operators. Chulaevsky and Sinaǐ [10] originally established AL for two-frequency Schrödinger operator almost everywhere in phase space. For the frequency of more than two dimensional case, the same result was obtained in [4]. However, these results were established directly in infinite volume, whereas localization in finite volume was not studied. A major breakthrough was achieved by Goldstein, Schlag, and Voda [15, 16], who pioneered analytical techniques to establish both finite and full scale localization for MF-QP Schrödinger operators. Based on the tools obtained in [15], they further derived a surprising result that the spectrum of this kind of operators is actually a single interval. Here we would also like to mention the recent papers [28] and [29] on MF-QP operators.
Substantial progress has also been made for CMV matrices. Localization has been established for various driving dynamics, including skew-shifts [9, 21], single-frequency quasi-periodic potentials [27], random perturbations [30], and hyperbolic toral automorphisms [20]. Meanwhile, the spectral theory in the small quasi-periodic regime has been explored, revealing purely absolutely continuous spectrum [18] and Cantor spectrum [19] in different settings. However, the understanding of MF-QP CMV matrices remains notably limited.
@@
To our knowledge, the finite and full scale localization for MF-QP CMV matrices remains unaddressed. To do that, we need large deviation theorem (LDT) for the entries of the transfer matrix. Comparing with the results in [13], there are some difficulties we have to overcome. This is primarily due to the more complex structure of CMV matrices. It is known to all, Schrödinger operator is a real tri-diagonal matrix, while the CMV matrix is a complex five-diagonal matrix, then the analysis about the corresponding characteristic determinant (entries of the transfer matrix) is more complicated. To treat this issue, we decompose the matrix into two parts and consider two cases; see the proof of Lemma 3.15 in detail. Consequently, the large deviation theorem for the entries of the transfer matrix follows.
@@
In addition, considering the restriction of the extended CMV matrix to a finite interval, from [17] and [30], the Poisson formula of CMV matrices depends on the parity of the endpoints of this interval. This also makes the analysis of the covering form of LDT more difficult to some extent. Thus we consider a subinterval of the above finite interval instead, then we get the covering lemma. Since the subinterval is a closed interval, the covering lemma can be applied to the original interval. Then the covering form of LDT can be obtained.
@@
Furthermore, the restriction of the Schrödinger operator to a finite interval results in a Hermitian matrix with many favorable properties. In contrast, the restriction of the CMV matrix results in a unitary matrix after modifying the boundary conditions, and its properties are less favorable than those of a Hermitian matrix. For example, there is a famous Cauchy Interlacing Theorem for the eigenvalues of Hermitian matrices. But we can not expect such a result for unitary matrices. This directly leads to a challenge in estimating the number of zeros for the characteristic determinant within a disk. Thus, the factorization of the characteristic determinant becomes more complex.
Furthermore, eliminating double resonances in the multi-frequency setting requires a more intricate scheme than in the single-frequency case [27]. Instead of removing a set of frequencies for a fixed phase, we must independently control and eliminate two separate sets of bad frequencies and phases, a process that is inherently more complex.
@@
The rest of this paper is organized as follows. We describe the settings in Section 2. In Section 3, we give the estimates of the Lyapunov exponent and prove LDT for the transfer matrix and the entries of the transfer matrix respectively. Furthermore, with the aid of Cartan’s estimate and Poisson formula, we get the spectral and covering form of LDT. Finally, we prove the finite scale localization of the MF-QP CMV matrices, see Theorem 3.26. In Section 4, we derive a result on elimination of double resonances by applying the theory of semialgebraic sets. In Section 5, we study the elimination of double resonances between no double resonances intervals. Based on the analysis in previous sections, we obtain the full scale localization of the MF-QP CMV matrices in Section 6, see Theorem 6.7. We put some useful definitions and lemmas in the appendix section, that is, Section 7.
@@ 2 Preliminaries
Let $\mathbb{D}=\{z:|z|<1\}$ denote the open unit disk in $\mathbb{C}$, and let $\mu$ be a nontrivial probability measure on $\partial\mathbb{D}=\{z:|z|=1\}$. Consequently, the functions $1,z,z^{2},\ldots$ are linearly independent in the Hilbert space $L^{2}(\partial\mathbb{D},d\mu)$. Let $\Phi_{n}(z)$ be the monic orthogonal polynomials, that is, $\Phi_{n}(z)=P_{n}[z^{n}]$, $P_{n}\equiv$ projection onto $\{1,z,z^{2},\ldots,z^{n-1}\}^{\bot}$ in $L^{2}(\partial\mathbb{D},d\mu)$. The orthonormal polynomials are then given by
@@
The Verblunsky coefficients $\{\alpha_{n}\}_{n=0}^{\infty}$ ($\alpha_{n}\in\mathbb{D}$) obey the equation
@@
Applying the Gram-Schmidt procedure to $\{1,z,z^{-1},z^{2},z^{-2},\cdots\}$, we obtain a CMV basis $\{\chi_{j}\}_{j=0}^{\infty}$, and the matrix representation of multiplication by $z$ relative to the CMV basis gives rise to the CMV matrix $\mathcal{C}$,
@@
$\mathcal{C}=\left(\begin{array}[]{ccccccc}\overline{\alpha}_{0}&\overline{\alpha}_{1}\rho_{0}&\rho_{1}\rho_{0}&&&&\\ \rho_{0}&-\overline{\alpha}_{1}\alpha_{0}&-\rho_{1}\alpha_{0}&&&&\\ &\overline{\alpha}_{2}\rho_{1}&-\overline{\alpha}_{2}\alpha_{1}&\overline{\alpha}_{3}\rho_{2}&\rho_{3}\rho_{2}&&\\ &\rho_{2}\rho_{1}&-\rho_{2}\alpha_{1}&-\overline{\alpha}_{3}\alpha_{2}&-\rho_{3}\alpha_{2}&&\\ &&&\overline{\alpha}_{4}\rho_{3}&-\overline{\alpha}_{4}\alpha_{3}&\overline{\alpha}_{5}\rho_{4}&\\ &&&\rho_{4}\rho_{3}&-\rho_{4}\alpha_{3}&-\overline{\alpha}_{5}\alpha_{4}&\\ &&&&\ddots&\ddots&\ddots\end{array}\right),$
@@
$\mathcal{E}=\left(\begin{array}[]{cccccccc}\ddots&\ddots&\ddots&&&&&\\ &-\overline{\alpha}_{0}\alpha_{-1}&\overline{\alpha}_{1}\rho_{0}&\rho_{1}\rho_{0}&&&&\\ &-\rho_{0}\alpha_{-1}&-\overline{\alpha}_{1}\alpha_{0}&-\rho_{1}\alpha_{0}&&&&\\ &&\overline{\alpha}_{2}\rho_{1}&-\overline{\alpha}_{2}\alpha_{1}&\overline{\alpha}_{3}\rho_{2}&\rho_{3}\rho_{2}&&\\ &&\rho_{2}\rho_{1}&-\rho_{2}\alpha_{1}&-\overline{\alpha}_{3}\alpha_{2}&-\rho_{3}\alpha_{2}&&\\ &&&&\overline{\alpha}_{4}\rho_{3}&-\overline{\alpha}_{4}\alpha_{3}&\overline{\alpha}_{5}\rho_{4}&\\ &&&&\rho_{4}\rho_{3}&-\rho_{4}\alpha_{3}&-\overline{\alpha}_{5}\alpha_{4}&\\ &&&&&\ddots&\ddots&\ddots\end{array}\right).$
In this paper, we consider a sequence of Verblunsky coefficients generated by an analytic function $\alpha(\cdot):\mathbb{T}^{d}\rightarrow\mathbb{D}$, $\alpha_{n}(x)=\alpha(T^{n}x)=\alpha(x+n\omega)$, $d>1$. Here, $T$ is the shift map defined by $x=x+\omega, 𝕋:=/\mathbb{T}$$\mathbb{R}$$\mathbb{Z}$, and $x,\omega\in\mathbb{T}^{d}$ are referred to as phase and frequency respectively. Moreover, the frequency vector $\omega\in\mathbb{T}^{d}$ satisfies the standard Diophantine condition
@@
for all nonzero $k\in\mathbb{Z}^{d}$, where $p>0$, $q>d$ are some constants. $\|\cdot\|$ denotes the distance to the nearest integer and $|\cdot|$ stands for the sup-norm on $\mathbb{Z}^{d}$, that is, $|k|=|k_{1}|+|k_{2}|+\cdots+|k_{d}|$, $k_{i}$ is the $i$-th element of the vector $k$. For the sake of convenience, we denote $\mathbb{T}^{d}(p,q)\subset\mathbb{T}^{d}$ be the set of $\omega$ satisfying (2.2).
@@
Taking the Szegő dual to both sides of (2.5), we obtain
@@
$\left(\begin{array}[]{c}\varphi_{n+1}\\ \varphi_{n+1}^{*}\end{array}\right)=S(\omega,z;x+n\omega)\left(\begin{array}[]{c}\varphi_{n}\\ \varphi_{n}^{*}\end{array}\right),$
@@
$S(\omega,z;x)=\frac{1}{\rho(x)}\left(\begin{array}[]{cc}z&-\overline{\alpha}(x)\\ -\alpha(x)z&1\end{array}\right).$
@@
$M(\omega,z;x)=\frac{1}{\rho(x)}\left(\begin{array}[]{cc}\sqrt{z}&-\frac{\overline{\alpha}(x)}{\sqrt{z}}\\ -\alpha(x)\sqrt{z}&\frac{1}{\sqrt{z}}\end{array}\right)\in\mathbb{S}\mathbb{U}(1,1),$
@@
$\lim_{n\rightarrow\infty}u_{n}(\omega,z;x)=\lim_{n\rightarrow\infty}L_{n}(\omega,z)=L(\omega,z)$ (2.12)
exists for a.e.xa.e.\;x. Note that
@@ 3.1 LDT Estimate for the Monodromy
[14, Proposition 9.1] Let $d$ be a positive integer. Suppose $u:D(0,2)^{d}\rightarrow[-1,1]$ is subharmonic in each variable; i.e., $z\rightarrow u(z_{1},z_{2},\ldots,z_{d})$ is subharmonic for any choice of $(z_{2},\ldots,z_{d})\in D(0,2)^{d-1}$ and similarly for each of the other variables. Assume furthermore that for some $n\geq 1$
@@
Here $\langle u\rangle=\int_{\mathbb{T}^{d}}u(\theta)d\theta$. If $d=2$ then the range is $0<\tau<\frac{1}{3}-\varepsilon_{2}$ and $\sigma=\frac{1}{3}-\tau-\varepsilon_{2}$ where $\varepsilon_{2}\rightarrow 0$ as $\varepsilon_{1}\rightarrow 0$.
Based on the above lemma, one can obtain the following LDT, a fundamental tool in localization theory.
@@
Under the framework of analytic continuation in several complex variables, we note that in Szegő cocycle map $\overline{\alpha}(x)=\overline{\alpha(\overline{x}_{1},\overline{x}_{2},\ldots,\overline{x}_{d})}$. Thus, $\overline{\alpha}(x)$ is analytic. Then $u_{n}$ is a continuous subharmonic function bounded by $1$ in $D(0,1)^{d}$. Moreover, $u_{n}$ satisfies the conditions in the above lemma, and (3.3) follows directly from (3.2). ∎